OpenStudy (anonymous):

3x+3y=6 whats the slope draw it

OpenStudy (anonymous):

Simplifying 3x + -3y = 6 Solving 3x + -3y = 6 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3y' to each side of the equation. 3x + -3y + 3y = 6 + 3y Combine like terms: -3y + 3y = 0 3x + 0 = 6 + 3y 3x = 6 + 3y Divide each side by '3'. x = 2 + y Simplifying x = 2 + y

OpenStudy (anonymous):

I'm not so sure how you would draw the slope but here's and example:Example 14: Write the Slope-Intercept Equation of a line through point (2, 5). The new line is to be parallel to the line 9x - 3y = -6. Let's find the slope of the line 9x - 3y = -6 by changing its equation to Slope Intercept form.

OpenStudy (anonymous):

- 3y = -9x - 6 y = 3x + 2 We can see that the slope of the line is m = 3. Since the new line is to be parallel to 9x - 3y = -6 , it also has m = 3. However, the new line is to go through the point (2, 5).

OpenStudy (anonymous):

Given the slope of a line and one point on the line, we can now use the Point-Slope Equation to find the equation of the requested line. y - 5 = 3 (x - 2) y - 5 = 3x - 6 and since we were asked to find the Slope-Intercept Equation, let's isolate the y-variable as follows y = 3x - 6 + 5 y = 3x - 1

OpenStudy (anonymous):

Hope I Helped! :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Your Welcome! :)